TurtleJonsi:

The length of a rectangle is 5 more than 3 times its width. Find the length and width of the rectangle with an area of 182 units. A. Length: 26, Width: 7 B. Length: 7, Width: 26 C. Length: 2, Width: 91 D. Length: 91, Width: 2

itsmehjay:

Dont just give out the answer, instead help them understand how to get the answer

Perla112236:

u have to dived to get the answer i think

TurtleJonsi:

Thank you!! 🖤

Vocaloid:

let L = length and W = width The length of a rectangle is 5 more than 3 times its width. ---> L = 3W + 5 area = L*W. if you substitute L = 3W + 5, you get area = (3W+5)(W) = 182 solve for W. since all the widths in the answer choices are different, you can technically stop here, but to get the length, go back to L = 3W + 5 and plug in your W value.

TurtleJonsi:

@Vocaloid Thank you! 🖤